A student has two loans totaled $45,000. The interest of one loan is 5.1% and the interest of the other loan is 3.8%. After one year, the student has paid $1953.10 in interest. What was the amount of each loan?

Answer by mananth(16946)

A student has two loans totaled $45,000. The interest of one loan is 5.1% and the interest of the other loan is 3.8%. After one year, the student has paid $1953.10 in interest. What was the amount of each loan?
...
total loan 45,000
5.1%-------------- $x
3.8%---------------$45000-x
..
Interest ----------$1953.10
..
5.1x+3.8(45000-x)=1953.10*100
5.1x-3.8x+171000=195310
1.3x=24310
/1.3
..
x= $18,700=======> @ 5.1%
45000-18700=26300
$26300========>@3.8%